# IntelliRadar: A Comprehensive Platform to Pinpoint Malicious Packages Information from Cyber Intelligence

### v4

@@ Abstract.
conference: International Conference on Software Engineering; 12-18 April, 2026; Rio De Janeiro, Brazilcopyright: none
@@ 1. Introduction
With the widespread global use of open-source software, malicious actors have discovered an effective means to disseminate malicious code through open-source platforms. Especially, TPL registries, such as NPM and PyPI, have become disaster areas of malicious code, in which attackers deliberately upload packages, embedded with malicious code, and induce downstream users to include them as dependencies. These malicious packages often carry viruses, Trojans, ransomware, etc. (npm_remote_trojans, ; npm_stole_pwd, ; pypi_malicious_trojans, ), and stealthily infiltrate user systems by masquerading as regular libraries or tools. Unlike unintentional vulnerabilities, these are deliberately created by attackers with malicious intent.
To address the threat of malicious packages, academic and industrial researchers have worked to prevent their spread by investigating malicious package types, taxonomies (guo2023empirical, ; zhou2024large, ; zhou2024oss, ), and attack surfaces (okafor2022sok, ; ladisa2023journey, ; ladisa2023sok, ), as well as developing detection tools (huang2024donapi, ; li2023malwukong, ; vu2023bad, ; zahan2024shifting, ); simultaneously, software composition analysis tools like Snyk (snykweb, ), BlackDuck (BlackDuckweb, ), OWASP Dependency Check (OWASPweb, ), and Dependabot (dependabotweb, ) with security databases are used to identify malicious dependencies, relying on platforms such as GitHub Advisory (advisory_database_ref, ), NVD (nvdweb, ), and OSV (osv_database_ref, ) for updates. However, despite the abundance of detection methods, the malicious packages discovered by these researchers are typically disclosed through scattered unstructured web pages, making it difficult for this information to effectively reach downstream users and developers. Consequently, many already identified malicious packages still remain on mirror servers. -up to 72.4% of malicious PyPI packages remain on mirror servers after being reported (guo2023empirical, ), with the colorwed package (snyk_colorwed, ) being downloaded over 1,000 times after identification as malicious, revealing both a lack of security awareness and increased risk as these packages become known to potential attackers during this period.
@@
To fill these gaps, we propose a comprehensive approach IntelliRadar to construct a platform to pinpoint the coverage, timeliness, and accuracy of automated collection and processing of malicious package intelligence. Specifically, for challenge 1), starting from existing malicious package reports, we conducted a thorough exploration of intelligence sources by summarizing domain-specific keywords and snowballing results from search engines, to identify as many intelligence sources as possible. For challenge 2), we incorporate the large language models (LLMs) to extract the key information of malicious packages precisely. Unlike general cyber threat intelligence approaches such as CTIKG (huang2024ctikg, ) and SecBERT (host2023constructing, ), our task requires distinguishing malicious package names from benign packages in the same text and performing ecosystem-specific semantic understanding. To handle potential LLM inaccuracy (i.e., hallucination (huang2023surveyhallucinationlargelanguage, )), we introduced a multi-faceted approach to prompt LLMs with domain knowledge of malicious packages. For challenge 3), we conduct an empirical analysis on conflicted information among intelligence, and based on their correctness, we introduce a voting mechanism based on recency to cross-validate intelligence from different sources, for the same malicious packages.
@@
$\bullet$ We constructed a comprehensive and human-validated dataset containing intelligence on 34,313 malicious package names, establishing the largest known database for PyPI and NPM package managers to date, which is publicly accessible through our website (website, ).
@@ 2. Motivation
$\bullet$ Delay in Intelligence Propagation In the propagation chain of malicious packages, significant delays exist between initial discovery and eventual inclusion in SSC security databases. Research by Jacobs et al.(10.1093/cybsec/tyaa015, ) indicates that attackers can exploit such information propagation delays to compromise downstream users’ software security. An typical motivating example is illustrated in Figure 1, the propagation timeline of the malicious PyPI package colorwed demonstrates this issue: while the package was discovered and reported on social media(colorwed_twitter, ) on the same day it was uploaded to the PyPI registry (December 23, 2022), and subsequently reported by multiple security companies (JFrog on November 24(colorwed_jfrog, ), Phylum and Sontype on December 15 (colorwed_phylum, ; colorwed_sonatype, )), it took approximately one month (December 21) before being included in software composition analysis tools like Snyk’s database. This case clearly demonstrates the critical timing deficiencies in current malicious package intelligence propagation mechanisms.
@@ 3.1. Intelligence Source Identification
To acquire the existing malicious package reports, we first collect the malicious package list (including 2,351 PyPI and 1,984 NPM packages) from the existing well-known dataset of malicious packages (ohm2020backstabber, ), and cross-map it with Snyk and OSV databases (osv_database_ref, ; snyk_database_ref, ), which are well verified with reference links, to collect all reference intelligence reports. Based on this corpus, we then select the most relevant keywords that are commonly used in these intelligence reports. Specifically, we collect two groups of keywords,
$\bullet$ Common Keywords, we aim to identify the keywords that are most commonly used in these malicious intelligence reports. To this end, we apply the Latent Dirichlet Allocation (LDA) model (tong2016text, ) to cluster the text topics from the existing reports. After excluding stopwords and irrelevant common words, we selected 17 common keywords that are related to malicious packages out of the top 100 clustered topics.
$\bullet$ Special Keywords, apart from these common keywords, there could also be certain words that are only common in specific report. To this end, we adopt the Term Frequency-Inverse Document Frequency (TF-IDF) algorithm (qaiser2018text, ) to identify the top 10 most frequent keywords that are not included in the common keywords in each report. After excluding irrelevant words, we aggregate these words as the special keywords. Interestingly, most of these selected special keywords are the names of malicious packages and their specific attributes, such as execution and control. The detailed list of keywords are presented in Table 1.
@@
Common Keywords package, security, malicious, attacker,
account, user, registry, code, software,
github, malware, repository, dataset,
infected, script, vulnerability, workflow
@@ 3.2. Intelligence Collection
After identifying intelligence sources, we collected all webpage content for historical intelligence extraction and continuous monitoring. As shown in Figure 3, we classified these sources into five functional categories. Each category employs distinct structures for SSC intelligence distribution: Cybersecurity companies utilize specialized second- or third-level domains, Security News sites implement keyword indexing systems, while developer communities, code platforms, and social networks distribute intelligence through specific accounts. These structural patterns enabled targeted collection from relevant subdomains and accounts, eliminating the need for comprehensive site-wide crawling. However, even within these targeted SSC-related intelligence sources, many webpages still contain content irrelevant to malicious package intelligence. To overcome this challenge, we established clear inclusion and exclusion criteria for filtering relevant webpages. Our filtering employs a two-tier approach: ecosystem-specific keywords (e.g., ”pypi,” ”npm”) as mandatory criteria, and common security keywords from  Table 1 (e.g., ”malicious,” ”security,” ”package”) as optional filters. This mechanism significantly reduces computational overhead while maintaining comprehensive coverage. To effectively collect webpage content from these Intelligence sources, we developed specialized extraction tools that parse HTML elements. Our parser targets information-rich elements including tables and lists, which frequently contain package details, as well as paragraphs (p), code snippets (code), and headings  (h1-h3) that often present Indicators of Compromise (IOCs) and technical analyses. Additionally, we extract data from iframes, particularly from sources that load CSV files containing detailed malicious package information.
@@ 3.3. Malicious Package Intelligence Extraction
As shown in Figure 4 (1), the web content often includes a significant amount of dispersed information related to malicious packages, which is highlighted in yellow for clarity and emphasis. To efficiently extract the key intelligence from these texts, we employed LLM to enhance the information extraction process (min2023recent, ; wang2023gpt, ). However, LLM still faces issues like hallucination, limiting extraction accuracy (ling2023improving, ). To address this, we first preprocess the web content to extract potential names of malicious packages. Then, we feed both the extracted potential malicious package information and the web content together into LLM to extract more detailed and accurate malicious package information (martino2023knowledge, ).
@@
The potential malicious package name extraction process is illustrated in Figure 4 (2). We have observed that malicious packages on PyPI and NPM often try to deceive downstream users into downloading them by creating package names that are similar to benign packages, such as through typosquatting (tahir2018s, ; vu2020typosquatting, ) creating package names with minor letter variations. As a result, the names of these malicious packages often aren’t real words found in English dictionaries. Based on this observation, we designed a two-step extraction pipeline. First, we employ the regular expression r”(?:@[a-zA-Z0-9-]+/)?[a-zA-Z0-9][a-zA-Z0-9._-]+” to extract strings matching package naming conventions from webpage content. This pattern captures NPM scoped packages (prefixed with @), NPM regular packages, and PyPI packages according to their standard naming specifications. Subsequently, we apply filtering operations to the extracted candidate package names: removing duplicates to eliminate redundancy, stripping trailing punctuation marks such as periods and commas, and filtering out common English vocabulary and stopwords, thereby retaining potential malicious package names.
@@
To effectively analyze malicious package intelligence, we employ CoT reasoning by decomposing the analysis into three components: entity extraction, entity relationship analysis, and entity verification. This decomposition naturally aligns with how security analysts process and validate intelligence data. Entity extraction forms the foundation by identifying discrete pieces of intelligence data, while relationship analysis reveals the broader patterns and connections within this data. The verification phase then ensures the integrity of both the extracted entities and their relationships.
@@
(1) Entity Extraction (Fig 4a): Identify and extract key entities related to malicious packages from the original text. The prompt design consists of six components: 1) task description about extracting malicious package related entities from the content, 2) the intelligence content, 3) 9 types of entities and their definitions including Package Name, Version, Date of Discovery, Repository URL, Method of Attack, Discoverer, Impacted Systems, Attack Vector, and Indicators of Compromise (IOCs), 4) potential package names identified in previous steps, 5) common malicious package naming patterns (typosquatting and misspelling), and 6) few-shot examples as reference. The LLM employs a Chain-of-Thought approach with progressive difficulty: first confirming malicious package names from potential candidates, then extracting directly observable information (versions, dates, URLs), followed by inferring complex semantic information requiring synthesis (attack methods, IOCs), and finally grouping packages with shared characteristics, ultimately producing structured JSON output containing the nine entity types for each identified malicious package.
@@
(3) Entity Verification (Fig 4c): This stage verifies the accuracy of the extracted information through LLM-based cross-validation and self-correction. The process inputs both the original intelligence content and the structured entities from the entity relationship analysis step. The LLM performs cross-validation by comparing the extracted entities and their relationships against the original text, identifying inconsistencies or errors, and making necessary corrections to ensure the final output maintains accuracy and completeness.
Entity Extraction Example. Table 2 demonstrates how IntelliRadar extracts malicious package intelligence entities from real unstructured web text. For the ”ef323refefeffe” package, our framework successfully extracts basic package information including name, version, and repository URL. More importantly, IntelliRadar can also infer malicious behaviors such as information stealing, Discord webhook abuse, and targeting Windows systems.
@@
Table 2. Entity Extraction: ”ef323refefeffe” Package
Entity Type Extracted Value
Package Name ef323refefeffe
Package Manager PyPI
Version 1.0
Discovery Date 2023-12-26
Repository URL https://socket.dev/pypi/package/ef323refefeffe
/files/1.0/tar-gz/ef323refefeffe-1.0/setup.py
Method of Attack stealing sensitive information,
abusing Discord webhooks
Discoverers Socket research team
Impacted Systems Windows
Attack Vector Exploiting Discord
IOC Indicators https://canvas.discord.com/api/webhooks/118942742960
1710100/JmLvp-Xz...a42Le_zNGMEa8p5V_VZxHh4VsEUpzVp6X
Collection Timestamp 2024-01-09
Note: Extracted from
https://socket.dev/blog/blank-grabber-python-package-steals-info-from-discord-and-telegram
3.4. Intelligence Aggretation
@@
Our analysis revealed that while most entity information is consistent across sources, conflicts do exist, particularly in attack methods and version numbers (5.13%), affected systems (2.45%), and repository URLs (0.41%). For these conflicting cases, we observed two important patterns: (1) later webpages tend to be more accurate as they often incorporate and reference earlier findings, and (2) information confirmed by multiple sources demonstrates higher reliability.
@@
Voting Mechanism:
  • For each package name $N$ of the malicious package information, vote on fields $V$, $R$, $M$, $D$, II, AA.
  • After excluding $NaN$ values, count the occurrences of each value. Let $E_{i}$ be the value of entity $E$ in the ith intelligence source. Define the voting count function $count(E_{i})$ to represent the number of occurrences of value $E_{j}$.
  • If there is a $\exists E_{i},E_{j}$ such that $count(E_{i})=count(E_{j})$, select the entry with the largest timestamp $T$: $E_{\text{final}}=\arg\max_{E_{i}}T_{i}$ For the discovery date $F$, select the earliest date: $F_{\text{final}}=\min_{i}F_{i}$ For IOCs $C$, merge all non-duplicate values: $C_{\text{final}}=\bigcup_{i}C_{i}$
@@
To ensure the accuracy of our final database, we implemented a two-step verification process. First, we verified package existence by validating against official npm and PyPI repository metadata APIs (npmregistry, ; pypistats, ; pypibigquery, ), which retain historical package information even after malicious packages are removed, ensuring extracted names correspond to actual packages rather than LLM hallucinations. Second, we confirmed maliciousness through multiple approaches: cross-referencing with established security databases (OSV, Snyk, GitHub Advisory), multi-source corroboration for packages reported by multiple independent sources, and manual verification for remaining packages, which proved efficient due to malicious packages typically being released in batches. Through this multi-layered verification approach, we established a reliable and accurate database of verified malicious packages.

### v6

@@ Abstract.
Supply Chain Security, Malicious Packages, Cyber Threat Intelligence, Third-Party Library
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773113isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software libraries and repositories
@@ 1. Introduction
With the widespread global use of open-source software, malicious actors have discovered an effective means to disseminate malicious code through open-source platforms. Especially, TPL registries, such as NPM and PyPI, have become disaster areas of malicious code, in which attackers deliberately upload packages, embedded with malicious code, and induce downstream users to include them as dependencies. These malicious packages often carry viruses, Trojans, ransomware, etc. (npm_remote_trojans; npm_stole_pwd; pypi_malicious_trojans), and stealthily infiltrate user systems by masquerading as regular libraries or tools. Unlike unintentional vulnerabilities, these are deliberately created by attackers with malicious intent.
To address the threat of malicious packages, academic and industrial researchers have worked to prevent their spread by investigating malicious package types, taxonomies (guo2023empirical; zhou2024large; zhou2024oss), and attack surfaces (okafor2022sok; ladisa2023journey; ladisa2023sok), as well as developing detection tools (huang2024donapi; li2023malwukong; vu2023bad; zahan2024shifting); simultaneously, software composition analysis tools like Snyk (snykweb), BlackDuck (BlackDuckweb), OWASP Dependency Check (OWASPweb), and Dependabot (dependabotweb) with security databases are used to identify malicious dependencies, relying on platforms such as GitHub Advisory (advisory_database_ref), NVD (nvdweb), and OSV (osv_database_ref) for updates. However, despite the abundance of detection methods, the malicious packages discovered by these researchers are typically disclosed through scattered unstructured web pages, making it difficult for this information to effectively reach downstream users and developers. Consequently, many identified malicious packages still remain on mirror servers. Up to 72.4% of malicious PyPI packages remain on mirror servers after being reported (guo2023empirical), with the colorwed package (snyk_colorwed) being downloaded over 1,000 times after identification as malicious, revealing both a lack of security awareness and increased risk as these packages become known to potential attackers during this period.
@@
To fill these gaps, we propose a comprehensive approach IntelliRadar to construct a platform to pinpoint the coverage, timeliness, and accuracy of automated collection and processing of malicious package intelligence. Specifically, for challenge 1), starting from existing malicious package reports, we conducted a thorough exploration of intelligence sources by summarizing domain-specific keywords and snowballing results from search engines, to identify as many intelligence sources as possible. For challenge 2), we incorporate the large language models (LLMs) to extract the key information of malicious packages precisely. Unlike general cyber threat intelligence approaches such as CTIKG (huang2024ctikg) and SecBERT (host2023constructing), our task requires distinguishing malicious package names from benign packages in the same text and performing ecosystem-specific semantic understanding. To handle potential LLM inaccuracy (i.e., hallucination (huang2023surveyhallucinationlargelanguage)), we introduced a multi-faceted approach to prompt LLMs with domain knowledge of malicious packages. For challenge 3), we conduct an empirical analysis on conflicted information among intelligence, and based on their correctness, we introduce a voting mechanism based on recency to cross-validate intelligence from different sources, for the same malicious packages.
@@
$\bullet$ We constructed a comprehensive and human-validated dataset containing intelligence on 34,313 malicious package names, establishing the largest database for PyPI and NPM package managers to date, which is publicly accessible through our website (website).
@@ 2. Motivation
$\bullet$ Delay in Intelligence Propagation In the propagation chain of malicious packages, significant delays exist between initial discovery and eventual inclusion in SSC security databases. Research by Jacobs et al.(10.1093/cybsec/tyaa015) indicates that attackers can exploit such information propagation delays to compromise downstream users’ software security. An typical motivating example is illustrated in Figure 1, the propagation timeline of the malicious PyPI package colorwed demonstrates this issue: while the package was discovered and reported on social media(colorwed_twitter) on the same day it was uploaded to the PyPI registry (December 23, 2022), and subsequently reported by multiple security companies (JFrog on November 24(colorwed_jfrog), Phylum and Sontype on December 15 (colorwed_phylum; colorwed_sonatype)), it took approximately one month (December 21) before being included in software composition analysis tools like Snyk’s database. This case clearly demonstrates the critical timing deficiencies in current malicious package intelligence propagation mechanisms.
@@ 3.1. Intelligence Source Identification
To acquire the existing malicious package reports, we first collect the malicious package list (including 2,351 PyPI and 1,984 NPM packages) from the existing well-known dataset of malicious packages (ohm2020backstabber), and cross-map it with Snyk and OSV databases (osv_database_ref; snyk_database_ref), which are well verified with reference links, to collect all reference intelligence reports. Based on this corpus, we then select the most relevant keywords that are commonly used in these intelligence reports. Specifically, we collect two groups of keywords,
$\bullet$ Common Keywords, we aim to identify the keywords that are most commonly used in these malicious intelligence reports. To this end, we apply the Latent Dirichlet Allocation (LDA) model (tong2016text) to cluster the text topics from the existing reports. After excluding stopwords and irrelevant common words, we selected 17 common keywords that are related to malicious packages out of the top 100 clustered topics.
$\bullet$ Special Keywords, apart from these common keywords, there could also be certain words that are only common in specific report. To this end, we adopt the Term Frequency-Inverse Document Frequency (TF-IDF) algorithm (qaiser2018text) to identify the top 10 most frequent keywords that are not included in the common keywords in each report. After excluding irrelevant words, we aggregate these words as the special keywords. Interestingly, most of these selected special keywords are the names of malicious packages and their specific attributes, such as execution and control. The detailed list of keywords are presented in Table 1.
@@
Common Keywords package, security, malicious, attacker, account, user, registry, code, software, github, malware, repository, dataset, infected, script, vulnerability, workflow
@@ 3.2. Intelligence Collection
After identifying intelligence sources, we collected all webpage content for historical intelligence extraction and continuous monitoring. As shown in Figure 3, we classified these sources into five functional categories. Each category employs distinct structures for SSC intelligence distribution: Cybersecurity companies utilize specialized second- or third-level domains, Security News sites implement keyword indexing systems, while developer communities, code platforms, and social networks distribute intelligence through specific accounts. These structural patterns enabled targeted collection from relevant subdomains and accounts, eliminating the need for comprehensive site-wide crawling. However, even within these targeted SSC-related intelligence sources, many webpages still contain content irrelevant to malicious package intelligence. To overcome this challenge, we established clear inclusion and exclusion criteria for filtering relevant webpages. Our filtering employs a two-tier approach: ecosystem-specific keywords (e.g., ”pypi,” ”npm”) as mandatory criteria, and common security keywords from  Table 1 (e.g., ”malicious,” ”security,” ”package”) as optional filters. This mechanism significantly reduces computational overhead while maintaining comprehensive coverage. To collect webpage content, we developed specialized HTML parsers targeting information-rich elements (tables, lists, paragraphs, code snippets, headings) and iframes containing malicious package details and IOCs.
@@ 3.3. Malicious Package Intelligence Extraction
As shown in Figure 4 (1), the web content often includes a significant amount of dispersed information related to malicious packages, which is highlighted in yellow for clarity and emphasis. To efficiently extract the key intelligence from these texts, we employed LLM to enhance the information extraction process (min2023recent; wang2023gpt). However, LLM still faces issues like hallucination, limiting extraction accuracy (ling2023improving). To address this, we first preprocess the web content to extract potential names of malicious packages. Then, we feed both the extracted potential malicious package information and the web content together into LLM to extract more detailed and accurate malicious package information (martino2023knowledge).
@@
The potential malicious package name extraction process is illustrated in Figure 4 (2). We have observed that malicious packages on PyPI and NPM often try to deceive downstream users into downloading them by creating package names that are similar to benign packages, such as through typosquatting (tahir2018s; vu2020typosquatting) creating package names with minor letter variations. As a result, the names of these malicious packages often aren’t real words found in English dictionaries. Based on this observation, we designed a two-step extraction pipeline. First, we employ the regular expression r”(?:@[a-zA-Z0-9-]+/)?[a-zA-Z0-9][a-zA-Z0-9._-]+” to extract strings matching package naming conventions from webpage content. Subsequently, we apply filtering operations to the extracted candidate package names: removing duplicates to eliminate redundancy, stripping trailing punctuation marks such as periods and commas, and filtering out common English vocabulary and stopwords, thereby retaining potential malicious package names.
@@
To analyze malicious package intelligence, we employ CoT reasoning with three components: entity extraction, relationship analysis, and verification. This decomposition aligns with how security analysts process intelligence: extracting discrete data, revealing patterns and connections, then ensuring integrity.
@@
(1) Entity Extraction (Fig 4a): Identify and extract key entities related to malicious packages. The prompt design consists of six components: 1) task description, 2) intelligence content, 3) 9 entity types including Package Name, Version, Date of Discovery, Repository URL, Method of Attack, Discoverer, Impacted Systems, Attack Vector, and IOCs, 4) potential package names from previous steps, 5) common malicious naming patterns (typosquatting and misspelling), and 6) few-shot examples. The LLM employs a Chain-of-Thought approach with progressive difficulty: first confirming malicious package names, then extracting observable information (versions, dates, URLs), followed by inferring complex semantics (attack methods, IOCs), and finally grouping packages with shared characteristics, producing structured JSON output containing the nine entity types.
@@
(3) Entity Verification (Fig 4c): This stage verifies the accuracy of the extracted information through LLM-based cross-validation and self-correction. The process inputs both the original intelligence content and the structured entities from the entity relationship analysis step. The LLM performs cross-validation by comparing the extracted entities and their relationships against the original text, identifying inconsistencies or errors, and making necessary corrections to ensure output maintains accuracy and completeness.
@@
3.4. Intelligence Aggregation
@@
Our analysis revealed that while most entity information is consistent across sources, conflicts do exist, particularly in attack methods and version numbers (5.13%), affected systems (2.45%), and repository URLs (0.41%). For these conflicting cases, we observed two patterns: (1) later webpages tend to be more accurate as they often incorporate and reference earlier findings, and (2) information confirmed by multiple sources demonstrates higher reliability.
@@
Voting Mechanism: For each package name $N$ of the malicious package information, vote on fields $V$, $R$, $M$, $D$, II, AA. After excluding $NaN$ values, count the occurrences of each value. Let $E_{i}$ be the value of entity $E$ in the ith intelligence source. Define the voting count function $count(E_{i})$ to represent the number of occurrences of value $E_{j}$. If there is a $\exists E_{i},E_{j}$ such that $count(E_{i})=count(E_{j})$, select the entry with the largest timestamp $T$: $E_{\text{final}}=\arg\max_{E_{i}}T_{i}$. For the discovery date $F$, select the earliest date: $F_{\text{final}}=\min_{i}F_{i}$. For IOCs $C$, merge all non-duplicate values: $C_{\text{final}}=\bigcup_{i}C_{i}$.
@@
To ensure the accuracy of our final database, we implemented a two-step verification process. First, we verified package existence by validating against official npm and PyPI repository metadata APIs (npmregistry; pypistats; pypibigquery), which retain historical package information even after malicious packages are removed, ensuring extracted names correspond to actual packages rather than LLM hallucinations. Second, we confirmed maliciousness through multiple approaches: cross-referencing with established security databases (OSV, Snyk, GitHub Advisory), multi-source corroboration for packages reported by multiple independent sources, and manual verification for remaining packages, which proved efficient due to malicious packages typically being released in batches. Through this multi-layered verification approach, we established a reliable and accurate database of verified malicious packages.